OpenStudy (dr.gray):

Which of the following is not a perfect square? A. 50 B. 49 C. 121 D. 1

OpenStudy (benji):

Do you know what a perfect square is? If not, what do you think it could mean?

OpenStudy (kittiwitti1):

http://www.mathwarehouse.com/arithmetic/numbers/what-is-a-perfect-square.php tl;dr A perfect square means the number \(A\) is a product of two values that are exactly the same, or \(squared\):\[\Large{b(b)=A\rightarrow b^{2}=A}\]

OpenStudy (kittiwitti1):

This also means that:\[\Large{\sqrt{A}=b}\]A "square root" symbol, or \(\sqrt{}\), means to reverse the squaring process.\[\large{b^{2}=A,\rightarrow\sqrt{A}=b}\]

OpenStudy (kittiwitti1):

@Dr.gray How are you doing? Are you confused?

OpenStudy (dr.gray):

yes i'm confused????????????????

OpenStudy (kittiwitti1):

Okay, yes, you're confused. Where are you confused?

OpenStudy (benji):

When you think of a perfect square, think about a number that can be multiplied by itself. So a perfect square could be 1*1 = 1 and 2*2=4 and 3*3=9, in all these 1,4, and 9 are perfect squares. If you can think of the numbers given, which numbers are multiplied by themselves to get 50, 49, 121, and 1?

OpenStudy (3mar):

Let me help, please. What are the two identical numbers (integers) which their product is one of the choices you offer?

OpenStudy (kittiwitti1):

Hint: choice B) \(49\) What is the square root of \(49\)? It is \(7\) because \(7\times7=49\). Apply the same idea to the rest of the options and check which one does not give you a whole number as a result.

OpenStudy (kittiwitti1):

(B is not the answer, btw)

OpenStudy (dr.gray):

it's A

OpenStudy (kittiwitti1):

Yes (:

OpenStudy (kittiwitti1):

Good job!